OpenStudy (anonymous):

What is the answer to the inequality -3≤x-2<1 Please can I get a quick answer

OpenStudy (cntrl_alt_defeated):

21

OpenStudy (anonymous):

Can I ask how you got that please?

OpenStudy (cntrl_alt_defeated):

9+10=21 so -3x-2=6+1=21

OpenStudy (cntrl_alt_defeated):

i hope u know im lying

OpenStudy (anonymous):

I was skeptical and trying to look at it differently so thank you for telling me you were lying....

OpenStudy (anonymous):

Serious answer: \[ -3 \le x - 2 < 1 \implies -1 \le x < 3. \]

OpenStudy (anonymous):

Ok, thank you I see how you got that. :)